A random sample of 11 observations was taken from normal population. The sample mean and standard deviation are 74.5 and 9 accordingly. Can we infer at 5% significance level that the population mean is greater than 70?


1
Expert's answer
2022-04-19T15:53:32-0400

"Let~\overline{x}=74.5,~\sigma=9,~n=11\\
\mu=70,~\alpha=0.05"

"z=\frac{\overline{x}-\mu}{\sigma}\sqrt{n}=\frac{74.5-70}{9}\sqrt{11}\approx1.658"

"P(z>1.658)=1-P(z<1.658)\\
P(z<1.658)>P(z<1.65)"

From z-score table: "P(z<1.65)=0.9505"

"P(z<1.658)>0.9505\\
P(z>1.658)<1-0.9505=0.0495<0.05=\alpha"

Since "P(z>1.658)<\alpha," we can infer that the population mean is greater than 70.